### v2

# : A Place is Worth a Bag of Learnable Queries

Abstract
In visual place recognition, accurately identifying and matching images of locations under varying environmental conditions and viewpoints remains a significant challenge. In this paper, we introduce a new technique, called Bag-of-Queries (BoQ), which learns a set of global queries, designed to capture universal place-specific attributes. Unlike existing techniques that employ self-attention and generate the queries directly from the input, BoQ employ distinct learnable global queries, which probe the input features via cross-attention, ensuring consistent information aggregation. In addition, this technique provides an interpretable attention mechanism and integrates with both CNN and Vision Transformer backbones. The performance of BoQ is demonstrated through extensive experiments on large-scale benchmarks. It consistently outperforms current state-of-the-art techniques including NetVLAD, MixVPR and EigenPlaces. Moreover, despite being a global retrieval technique (one-stage), BoQ surpasses two-stage retrieval methods, such as Patch-NetVLAD, TransVPR and R2Former, all while being orders of magnitude faster and more efficient. The code and model weights are publicly available at https://github.com/amaralibey/Bag-of-Queries.
1 Introduction
Visual Place Recognition (VPR) consists of determining the geographical location of a place depicted in a given image, by comparing its visual features to a database of previously visited places. The dynamic and ever-changing nature of real-world environments pose significant challenges for VPRĀ [35, 60]. Factors such as varying lighting conditions, seasonal changes and the presence of dynamic elements such as vehicles and pedestrians introduce considerable variability into the appearance of a place. Additionally, changes in viewpoint and image scale can expose previously obscured areas, further complicating the recognition process. These challenges are exacerbated by the operational constraints of VPR systems, which often need to operate in real-time and under limited memory. Consequently, there is a compelling need for efficient algorithms capable of generating compact yet robust representations.
With the rise of deep learning, numerous VPR-specific neural networks have been proposedĀ [6, 30, 43, 32, 22, 3, 2, 11, 4], leveraging Convolutional Neural Networks (CNNs) to extract high-level features, followed by aggregation layers that consolidate these features into a single global descriptor. Such end-to-end trainable architectures have been instrumental in enhancing the efficiency and performance of VPR systems.
Recently, Vision Transformers (ViT)Ā [20] have demonstrated remarkable performance in various computer vision tasks, including image classificationĀ [15], object detectionĀ [14, 33] and semantic segmentationĀ [61]. Their success can be attributed to their self-attention mechanism, which captures global dependencies between distant parts of the imageĀ [23]. However, current Transformer-based VPR techniquesĀ [49, 62, 57, 52], often rely on rerankingĀ [8], a post-processing step aimed at refining the initial set of candidate locations identified through global descriptor search. The reranking process is usually done with geometric verification (e.g. RANSAC) on stored local patch tokens, which is computationally and memory intensive. Moreover, the global retrieval process in Transformer-based methods, whether through specific aggregation of local patchesĀ [57, 49] or using the class token of the ViTĀ [62], has yet to reach the performance levels of non-Transformer-based approaches like MixVPRĀ [4] and CosPlaceĀ [61].
In this paper, we bridge this performance gap, by introducing a novel Transformer-based aggregation technique, called Bag-of-Queries (BoQ), that learns a set of embeddings (global queries) and employs a cross-attention mechanism to probe local features coming from the backbone network. This approach enables each global query to consistently seek relevant information uniformly across input images. This is in contrast with self-attention-based techniquesĀ [49, 57], where the queries are dynamically derived from the input itself. Furthermore, BoQ is designed for end-to-end training, thus seamlessly integrating with both conventional CNN and ViT backbones. Its effectiveness is validated through extensive experiments on multiple large-scale benchmarks, consistently outperforming state-of-the-art techniques, including MixVPRĀ [4], EigenPlacesĀ [12], and NetVLADĀ [6]. More importantly, BoQ, a single-stage (global) retrieval method that does not employ reranking, outperforms two-stage retrieval methods like TransVPRĀ [49] and R2FormerĀ [62], while being orders of magnitude more efficient in terms of computational and memory resources.
2 Related Works
Early visual place recognition (VPR) techniques relied on hand-crafted local features such as SIFTĀ [34], SURFĀ [9] and ORBĀ [42], which were aggregated into global descriptors using techniques like Bag-of-Words (BoW)Ā [40, 46, 21] or Vector of Locally Aggregated Descriptors (VLAD)Ā [28, 5]. BoW involves learning a visual vocabulary (or set of clusters), where each visual word (or cluster) represents a specific visual characteristic. Images are then represented as histograms, with bins indicating the frequency of each visual word. VLAD extends on this by capturing first-order statistics by accumulating the differences between local descriptors and their respective cluster centers.
CNN-based VPR. The advent of deep learning marked a significant shift in VPR techniques, with various aggregation architectures proposed. ArandjelovicĀ et al.Ā [6] introduced NetVLAD, a trainable version of VLAD that integrates with CNN backbones, achieving superior performance over traditional methods. Following its success, many researchers proposed variants such as CRNĀ [30], SPE-VLADĀ [55], GatedĀ NetVLADĀ [58] and SSL-VLADĀ [37]. Other approaches focus on regions of interests within feature mapsĀ [7, 45, 59]. Another key aggregation method in image retrieval is the Generalized Mean (GeM)Ā [41], which is a learnable form of global pooling. Building on GeM, BertonĀ et al.Ā [61] have recently introduced CosPlace, enhancing GeM by integrating it with a linear projection layer, achieving remarkable performance for VPR. Ali-beyĀ et al.Ā [4] introduced MixVPR, an all-MLP aggregation technique that achieved state-of-the-art performance on multiple benchmarks. Another facet of VPR is the training procedure, where the focus is on the loss functionĀ [3, 2, 12, 31].
Transformer-based VPR. The Transformer architecture was initially introduced for natural language processingĀ [48], and later adapted to Vision Transformers (ViT) for computer vision applicationsĀ [20]. In VPR, they have recently been used as backbone combined with MixVPRĀ [27, 26] or NetVLADĀ [51] resulting in a performance boost compared to CNN backbones. Furthermore, AnyLocĀ [29] used features extracted from a foundation model, called DINOv2Ā [38], combined with unsupervised aggregation methods such as VLAD, resulting in notable performance on a wide range of benchmarks.
Recently, two-stage retrieval strategy has gained prominence in Transformers-based VPR. It starts with global retrieval using global descriptors to identify top candidates, followed by a computationally-intensive reranking phase that refines the results based on local features. In this context, WangĀ et al.Ā [49] introduced TransVPR, which uses CNNs for feature extraction and Transformers for attention fusion to create global image descriptors, with additional patch-level descriptors for geometric verification. On the other hand, ZhangĀ et al.Ā [57] proposed ETR, a Transformer-based reranking technique that uses a CNN backbone for local and global descriptors. Their method leverages cross-attention between the local features of two images to cast the reranking as classification. Recently, ZhuĀ et al.Ā [62] proposed a unified framework integrating global retrieval and reranking, solely using transformers. For global retrieval, it employs the class token and utilizes other image tokens as local features for the reranking module. These two-stage techniques showed great performance, but at the expense of more computation and memory overheads. However, their global retrieval performance is still very limited compared to one-stage methods such as MixVPRĀ [4] and CosPlaceĀ [11]. In this work, we propose, Bag-of-Queries (BoQ), a Transformer-based aggregation technique for global retrieval, that outperforms existing state-of-the-art without relying on reranking. This makes BoQ particularly suitable for applications where computational resources are limited, yet high accuracy and efficiency are essential.
3 Methodology
In visual place recognition, effective aggregation of local features is crucial for accurate and fast global retrieval. To address this, we propose the Bag-of-Queries (BoQ) technique, a novel aggregation architecture that is end-to-end trainable and surprisingly simple, as depicted in Fig.Ā 2.
Our technique rely on Multi-Head Attention (MHA) mechanismĀ [48], which takes three inputs, queries (), keys () and values (), and linearly project them into multiple parallel heads. The output, for each head, is computed as follows:
In this mechanism, the queries play a crucial role. They act as a set of filters, determining which parts of the input (represented by keys and values ) are most relevant. The attention scores, derived from the dot product between queries and keys, essentially indicate to the model the degree of attention to give to each part of the input. We refer to self-attention when , and are derived from the same input, e.g. MHA(, , ). In contrast, cross-attention is the scenario where the query comes from a different source than the key and value.
Given an input image , we first process it through a backbone network, typically a pre-trained Convolutional Neural Network (CNN) or Vision Transformer (ViT), to extract its high-level features. For CNN backbones, we apply a convolution to reduce their dimensionality, whereas, in the case of a ViT backbone, we apply a linear projection for the same purpose. We regard the result as a sequence of local features of dimension , such as: . We then process through a sequence of Transformer-Encoder units and BoQ blocks. Each encoder transforms its input features, and passes the result to its subsequent BoQ block as follows:
Here, denotes the feature input to the -th Encoder unit, and represents the transformed output, which becomes the input for the next block in the pipeline.
[…]
Before using to aggregate information in the BoQ block, we first apply self-attention between them:
The self-attention operation allows the learnable queries to integrate their shared information during the training phase. Next, we apply a cross-attention between and and the input features of the current BoQ block:
This allows the learnable queries to dynamically assess the importance of each input feature by computing the relevance scores (attention weights) between and and aggregate them into the output .
Following this, the outputs from each BoQ block are concatenated to form a single output.
This ensures that the final descriptor combines information aggregated at different levels of the network, forming a rich and hierarchical representation. The final step involves reducing the dimensionality of the output using one or two successive linear projections with weights and , similar to the approach inĀ [3, 4].
This is followed by -normalization, to bring the global descriptor to the unit hypersphere, which optimizes the similarity search. The training is performed using pair-based (or triplet-based) loss functions that are widely used in the literatureĀ [3].
Relations to other methods. The Bag-of-Queries (BoQ) approach bears conceptual resemblance to Detection Transformer (DETR) modelĀ [14] in its use of object queries. However, BoQ is fundamentally different, as its global queries are exclusively used to extract and aggregate local feature information from the input; these queries do not contribute directly to the final representation, as demonstrated by the absence of any residual connection between the global queries and the output of the cross-attention. This is in contrast to DETR, where the object queries are integral part to the output, upon which object detection and classification is done.
In comparison to NetVLADĀ [6], which employs a set of learned cluster centers, and aggregates local features by calculating the weighted sum of their distances to each center. BoQ, leverages cross-attention between its learned queries and the input local feature to dynamically assess their relevance.
4 Experiments
In this section, we conduct comprehensive experiments to demonstrate the effectiveness of our technique, Bag-of-Queries (BoQ), compared to current state-of-the-art methods. First, we describe the datasets used, and then the implementation details. This is followed by the evaluation metrics. We then provide a detailed comparative analysis of performance, ablation studies and qualitative results.
4.1 Datasets
Our experiments are performed on existing benchmarks that present a wide range of real-world challenges for VPR systems. Tab.Ā 1 provides a brief summary of these benchmarks. MapillarySLS (MSLS)Ā [53] was collected using dashcams and features a broad range of viewpoints and lighting changes, testing the systemās adaptability to typical variations in ground-level VPR. Pitts250kĀ [46] and Pitts30k datasets, extracted from Google Street View, exhibit significant changes in viewpoint, which tests the systemās ability to maintain recognition across varying angles and perspectives. AmsterTimeĀ [54] presents a unique challenge with historical grayscale images as queries and contemporary color images as references, covering temporal variations over several decades. EynshamĀ [18] consists entirely of grayscale images, adding complexity due to the absence of color information. NordlandĀ [44] dataset was collected over four seasons, using a camera mounted on the front of a train. It encompasses scenes ranging from snowy winters to sunny summers, presenting the challenge of extreme appearance changes due to seasonal variations. Note that two variants of Nordland are used in VPR benchmarking: one uses a subset of 2760 summer images as queries against the entire winter sequence as reference images (marked withĀ *), while the other uses the entire winter sequence as query images against the entire summer sequence as reference images (marked withĀ **). SVOXĀ [10] stands out with its comprehensive coverage of weather conditions, including overcast, rainy, snowy, and sunny images, to test adaptability to meteorological changes. Additionally, SVOX Night subset focuses on nocturnal scenes, a challenging scenario for most VPR systems.
| Variations | |||||
| Dataset name | # quer. | # ref. | Viewpoint | Season | Day/Night |
| MSLSĀ [53] | k | ||||
| Pitts250kĀ [46] | k | k | |||
| Pitts30kĀ [46] | k | k | |||
| AmsterTimeĀ [54] | |||||
| EynshamĀ [18] | k | k | |||
| Nordland*Ā [56] | k | ||||
| Nordland**Ā [44] | k | k | |||
| St-LuciaĀ [36] | |||||
| SVOXĀ [10] | k | k | |||
| SPEDĀ [56] | |||||
4.2 Implementation details
Architecture. For our experiments, we primarily employ pre-trained ResNet-50 backbones for feature extractionĀ [25]. Importantly, our proposed Bag-of-Queries (BoQ) technique is architecturally agnostic and seamlessly integrates with various CNNs or Vision Transformer (ViT) backbones. When employing a ResNet, we crop it at the second last residual to preserve a higher resolution of local features. Additionally, we freeze all but the last residual block and add a convolution to reduce the number of channels, which decreases the computational and memory footprint of the self-attention mechanism. Implementing the BoQ model is straightforward with popular deep learning frameworks such as PyTorchĀ [39] and TensorFlowĀ [1], both providing highly optimized implementations of self-attention and cross-attention, which are the basic blocks of our technique.
Training. We train BoQ following the standard framework of GSV-CitiesĀ [3], which provides a highly accurate dataset of k images depicting k places. For the loss function, we use Multi-Similarity lossĀ [50], as it has been shown to perform best for visual place recognitionĀ [3]. We use batches containing places, each depicted by images resulting in mini-batches of images. We use the AdamW optimizer with a learning rate of depending the batch-size, and a weight decay of . The learning rate is multiplied by after each epochs. Finally, we train for a maximum of epochs (including epochs of linear warmup) using images resized to . Note that while current techniques such as NetVLADĀ [6], CosPlaceĀ [61] and EigenPlacesĀ [12] train on images of size āeffectively triple the pixel count of our chosen resolutionāour BoQ model still achieves better performance, underscoring its effectiveness.
Evaluation metrics. We follow the same evaluation metric of existing literature [6, 30, 53, 56, 49, 11, 4, 3, 2, 12], where the recall@ is measured. Recall@ is defined as the percentage of query images for which at least one of the top- predicted reference images falls within a predefined threshold distance, which is typically meters. However, for the Nordland benchmark, which features aligned sequences, the threshold is frame counts for Nordland** and only frame for Nordland*.
4.3 Comparison with previous works
| Method | Dim. | Pitts250k-test | MSLS-val | SPED | Nordland* | ||||||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| AVGĀ [6] | |||||||||||||
| GeM [41] | |||||||||||||
| NetVLAD [6] | |||||||||||||
| SPE-NetVLAD [55] | |||||||||||||
| Gated NetVLAD [58] | |||||||||||||
| Conv-AP [3] | |||||||||||||
| CosPlace [11] | |||||||||||||
| MixVPR [4] | |||||||||||||
| EigenPlaces [12] | |||||||||||||
| Multi-view datasets | Frontal-view datasets | ||||||||||||||||||||
| Method | Backbone | AmsterTime | Eynsham | Pitts30k | Nordland** | St Lucia |
| NetVLAD [6] | VGG-16 | ||||||||||||||||||||
| SFRS [22] | VGG-16 | ||||||||||||||||||||
| CosPlace [11] | VGG-16 | ||||||||||||||||||||
| EigenPlaces [12] | VGG-16 | ||||||||||||||||||||
[…]
| CosPlace [11] | ResNet-50 | ||||||||||||||||||||
[…]
| Extrac. (ms) | Rerank. (ms) | MSLS-val | Pitts30k | |||
| R@1 | R@5 | R@1 | R@5 | |||
| SP-SuperGlue [19] | ||||||
| DELG [13] | ||||||
| Patch-NetVLAD [24] | ||||||
| TransVPR [49] | ||||||
| R2Former [62] | ||||||
In this section, we present an extensive comparison of our proposed method, Bag-of-Queries (BoQ), with a wide range of existing state-of-the-art VPR aggregation techniques. This includes global average pooling (AVG)Ā [6], Generalized MeanĀ (GeM)Ā [41], NetVLADĀ [6] alongside its recent variants, SPE-NetVLADĀ [55] and GatedĀ NetVLADĀ [58]. Importantly, we compare against recent cutting-edge techniques, such as ConvAPĀ [3], CosPlaceĀ [61], EigenPlacesĀ [12] and MixVPRĀ [4], which currently hold the best scores in most existing benchmarks, setting a high standard for BoQ to measure up against in the field of VPR.
Moreover, despite BoQ being a global retrieval technique which does not employ reranking, we compare it with current state-of-the-art two-stage retrieval techniquesĀ [19, 13, 24, 49, 62]. These methods leverage geometric verification to enhance recall@k performance at the expense of increased computation and memory overhead.
Results analysis. In Tab.Ā 2, we present a comparative analysis across several challenging datasets, focusing on Recall@1 (R@1), Recall@5 (R@5), and Recall@10 (R@10). We also offer insight into the performance of each method with regard to the top-ranked retrieved images.
ā¢
On Pitts250k-test, BoQ outperforms all other methods with a marginal yet notable improvement as it is the first global technique that reaches the R@1 threshold on this benchmark.
ā¢
On MSLS-val and SPED, BoQ outperforms the second best methods (EigenPlaces and MixVPR) by and percentage points respectively.
ā¢
On the challenging Nordland* benchmark, known for its extreme seasonal variations, BoQ significantly outperforms all compared methods, indicating its robustness to severe environmental changes. Surpassing CosPlace and Eigenplaces by , and MixVPR by percentage points, which translates to over additional accurate retrievals within the top- results.
In Tab.Ā 3, we follow the same evaluation style proposed by BertonĀ et al.Ā [12], where benchmarks are categorized into multi-view datasets, where images are oriented in various angles, and frontal-view dataset where images are mostly forward facing. Top-1 retrieval performance (R@1) is reported for each technique.
ā¢
For the multi-view benchmarks, BoQ achieves highest R@1 on AmsterTime and Eynsham, and second best compared to EigenPlaces on Pitts30k-test. Both techniques show robustness to viewpoint changes in urban environments, while facing challenges with the AmsterTime dataset, which includes decades-old historical imagery.
ā¢
On frontal-view benchmarks, particularly on Nordland** (which comprises k queries), BoQ achieves R@1 of , which is and percentage points more than MixVPR, and CosPlace respectively. This translates to BoQ correctly retrieving an additional and images within the top-1 results.
ā¢
On SVOX, BoQ achieves best results under all conditions, especially on SVOX night, where it achieves R@1, outperforming the second best method (MixVPR) by percentage points. This highlights BoQās potential under low-light conditions.
Furthermore, we compare BoQ to two-stage retrieval techniques in Tab.Ā 4, by reporting performance on MSLS-val and Pitts30k-test. Although our technique does not perform reranking, its global retrieval performance surpasses that of existing two-stage techniques, including Patch-NetVLAD (2021), TransVPR (2022) and R2Former (2023). Importantly, this makes our approach more efficient in terms of memory and computation: over faster retrieval time compared to R2Former.
4.4 Ablation studies
Number of learnable queries. We conduct an ablation study to investigate the impact of varying the number of learnable queries, , on the performance of BoQ. We use a ResNet-50 backbone and two BoQ blocks. The results are presented in Tab.Ā 5. We observe that performance improves as the number of queries increases. However, the benefits of additional queries vary across datasets. In less diverse environments, such as Pitts30k, adding more queries yields marginal performance improvements (increasing from to queries results in a mere R@1 performance gain). In contrast, the results on AmsterTime, which features highly diverse images spanning decades, demonstrate that the model benefits significantly from additional queries. This aligns with the underlying intuition of BoQ, where each query implicitly learns a distinct universal pattern.
| Size of | MSLS-val | Pitts30k-val | AmsterTime | |||
| R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | |
MSLS | Nordland | AmsterTime |
| MSLS-val | Pitts30k | Pitts250k | Nordland | |
| BoQ w/out self-att. | ||||
| BoQ with self-att. |
| Backbone | # params (M) | MSLS-val | SPED | Nordland* | |||
| R@1 | R@5 | R@1 | R@5 | R@1 | R@5 | ||
| ResNet-18 | |||||||
| ResNet-34 | |||||||
| ResNet-50 | |||||||
| ResNet-101 | |||||||
Number of BoQ blocks. To assess the influence of the number of BoQ blocks within our architecture, we conduct experiments by varying , the number BoQ blocks utilized. We use a ResNet-18 backbone followed by BoQ blocks, each comprising a bag of learnable queries. We report R@1 performance for each setup. The results, presented in Tab.Ā 6, demonstrate that even when paired with a lightweight backbone like ResNet-18, BoQ remains competitive against state-of-the-art methods such as CosPlaceĀ [11] and MixVPRĀ [4], which use ResNet-50 backbone. Best overall performance is observed when employing BoQ blocks.
Ablation on the self-attention. Applying self-attention on () brings more stability/performance by adding context between the queries. As shown in Tab.Ā 7, we train a BoQ model with ResNet-18 backbone for epochs. Adding self-attention between the learnable queries brings consistent performance improvement across various benchmarks. Note that the self-attentionās output can be cached during eval/test to avoid recomputation at every test iteration.
Backbone architecture. In Tab.Ā 8 we present a comparative performance of BoQ coupled with different ResNet backbone architectures. For each backbone, we crop it at the second last residual block and follow it with two BoQ blocks, each comprising learnable queries. The total number of parameters (trainable + non trainable) is provided. The empirical results showcase that when BoQ is coupled with ResNet-34, a relatively lightweight backbone, it achieves competitive performance compared to NetVLAD, CosPlace and EigenPlaces coupled with ResNet-50 backbone. Interestingly, using ResNet-101, a relatively deeper backbone, we did not achieve better performance than ResNet-50, which could be attributed to memory constraints necessitating a smaller training batch size.
4.5 Vision Transformer Backbones
All previous experiments in this paper were conducted using a ResNetĀ [25] backbone to ensure fairness against existing state-of-the-art methods. Recently, vision foundation models such as DINOv2Ā [38] have been introduced and rapidly adopted as the de facto backbone for various computer vision tasks. In this experiment, we evaluate the performance of BoQ when using a DINOv2 backbone. We employ the pre-trained base variant with 86M parameters, freezing all but the last two blocks to enable fine-tuning. We train on GSV-CitiesĀ [3] with images resized to and perform testing with images resized to . The results presented in Tab.Ā 9, demonstrate that DINOv2 further enhances the performance of BoQ, achieving new all-time high performance on all benchmarks.
| Test Dataset | DinoV2-BoQ | ResNet50-BoQ | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |
| MSLS-valĀ [53] | ||||||
| Pitts30k-testĀ [46] | ||||||
| Pitts250k-testĀ [46] | ||||||
| Nordland**Ā [44] | ||||||
| Nordland*Ā [56] | ||||||
| SPEDĀ [56] | ||||||
| AmesterTimeĀ [54] | ||||||
| EynshamĀ [18] | ||||||
| St-LuciaĀ [36] | ||||||
| SVOX-nightĀ [10] | ||||||
| SVOX-sunĀ [10] | ||||||
| SVOX-rainĀ [10] | ||||||
| SVOX-snowĀ [10] | ||||||
| SVOX-overcastĀ [10] | ||||||
| SVOX-allĀ [10] | ||||||
| Tokyo 24/7Ā [47] | ||||||
| MSLS-ChallengeĀ [53] | ā | ā | ā | |||
| SF-LandmarkĀ [16] | ā | ā | ā | |||
4.6 Visualization of the learned queries
[…]
Observing horizontally, the diversity in attention patterns across the learned queries becomes apparent. For instance, the first query (top row) appears to concentrate on fine-grained details within the feature maps, as indicated by the intense localized areas. In contrast, the second query (second row) shows a preference for larger regions in the input features, suggesting a more holistic capture of scene attributes.
5 Conclusion
In this work, we introduced Bag-of-Queries (BoQ), a novel aggregation technique for visual place recognition, which is based on the use of learnable global queries to probe local features via a cross-attention mechanism, allowing for robust and discriminative feature aggregation. Our extensive experiments on different large-scale benchmarks demonstrate that BoQ consistently outperforms current state-of-the-art techniques, particularly in handling complex variations in viewpoint, lighting, and seasonal changes. Furthermore, BoQ being a global (one-stage) retrieval technique, it outperforms existing two-stage retrieval methods that employ geometric verification for re-ranking, all while being orders of magnitude faster, setting a new standard for VPR research.
In future work, building upon the concepts discussed in Sec.Ā 3, the spatial information carried in the output of the last encoder, denoted as , presents an opportunity for further enhancement through the application of special reranking strategies.
Acknowledgement: This work has been supported by The Fonds de Recherche du QuƩbec Nature et technologies (FRQNT 254912). We gratefully acknowledge the support of NVIDIA Corporation with the donation of a Quadro RTX 8000 GPU used for our experiments.
Supplementary Material
6 More implementation details
Training image size. For our training implementation, we resized the images to pixels to maintain consistency with the training procedures of [4, 3, 2]. This choice of specific resolution directly influences the spatial resolution of the resulting feature maps, which is 20 when using a cropped ResNet backboneĀ [25]. This choice also allows for the use of larger batch sizes.
Inference image size. Considering that many benchmarks contain images of varying sizes and aspect ratiosāand often at higher resolutionsā we resize the images to a resolution slightly higher than pixels while preserving the original aspect ratio. This approach maintains the integrity of the scenes by keeping the original aspect ratio, and allows the learned queries to interact with bigger feature maps. In Tab.Ā 10 we perform testing at different image sizes (), using a BoQ model trained with images. As we can see, when the images are resized to a height of pixels, there is a consistent improvement in Recall@1 across almost all datasets. This experiment suggests that heights of and may represent an optimal balance between image detail and the modelās capacity to extract and utilize informative features. Note that the performance gains from resizing to these heights are marginal compared to the baseline size of pixels.
AmsterTime | Eynsham | St-Lucia | SVOX (all) | Nordland** | ||
7 Interpretability of the learned queries
In this section, we demonstrate how the learned queries in BoQ can be visually interpreted through their direct interactions with the feature maps via cross-attention mechanisms. To do so, we examine their behavior in images of the same location under viewpoint changes, occlusions, and varying weather conditions.
The cross-attention mechanisms in our BoQ model have been instrumental in achieving fine-grained feature discrimination, as demonstrated by Fig.Ā 5, Fig.Ā 6 and Fig.Ā 7. These figures provide a visualization of the learned queriesā attention patterns across diverse urban scenes and under various environmental conditions.
Fig.Ā 5 demonstrates the modelās temporal robustness, displaying consistent attention across images of the same location captured at different times. The learned queries reliably focus on distinctive features, such as buildings, foliage, and poles, despite variations in viewpoint, lighting, and weather conditions.
Moving objects within a scene often pose a challenge for VPR techniques. Nonetheless, as shown in Fig.Ā 6 our learned queries focus their attention towards static elements of the environment, avoiding moving objects like vehicles.
Fig.Ā 7 underscores the specialization of the learned queries, showcasing their selective focus on relevant features, such as vegetation and buildings. This selective attention is indicative of our modelās ability to interpret complex visual information within the environment.